What is (-2x+14)-15(200x-30)-(2x-10)= what
Answer by addingup(3677)

(-2x+14)-15(200x-30)-(2x-10)=
Let's consolidate the elements in the last parenthesis with those in the first parenthesis:
- -10= +10 and 10+14= 24
-2x-2x= -4x
So now we have:
24-4x-15(200x-30) = 0
Now distribute (multiply) -15 over 200x and -30:
-15(200x-30) = 450-3000x:
24-4x+450-3000x
Group like terms:
24+450-4x-3000x
Add/subtract:
474-3004x
474-3004x = 0
474= 0+3004x
474= 3004x
474/3004 = x
factors of 3004: 1, 2, 4,...
factors of 474: 1, 2, 3, 6,....
Divide by 2:
237/1502 No common factors, we can't simplify any more.